Using the fact that if z = r(cosθ +isinθ) then z2 = r2(cosθ +isinθ), find the square root of

2√3 -2i


1
Expert's answer
2021-09-15T03:05:29-0400

"\displaystyle
\textsf{Let}\,\,\, z = 2\sqrt{3} - 2i\\

\arg{z} = \arctan\left(\frac{-2}{2\sqrt{3}}\right) = -\frac{\pi}{6}\\

|z| = \sqrt{12 + 4} = 4\\

\implies z = 4e^{-\frac{i\pi}{6}}\\

\begin{aligned}
z^{\frac{1}{2}} &= \pm 2e^{-\frac{i\pi}{12}}
\\&= \pm 2\cos\left(\frac{\pi}{12}\right) - \left(\pm i2\sin\left(\frac{\pi}{12}\right)\right) \\&= \pm 2\cos\left(\frac{\pi}{12}\right) \mp i2\sin\left(\frac{\pi}{12}\right)
\end{aligned}\\
\textsf{It implies that the square root is}\\2\cos\left(\frac{\pi}{12}\right) - i2\sin\left(\frac{\pi}{12}\right) \,\,\textup{or}\\
-2\cos\left(\frac{\pi}{12}\right) + i2\sin\left(\frac{\pi}{12}\right)"